OpenStudy (souvik):

3l-2m+3n=0 l-2m-n=0 find l,m & n...

OpenStudy (anonymous):

3l - 2m + 3n = 0 l - 2m - n = 0 2l + 4n = 0 l = -2n -3n - 2m = 0 m = -2/3 n

OpenStudy (anonymous):

\[3l-2m+3n=0\]\[l-2m-n=0\] Subtracting both equations,\[2l+4n=0\]\[l=-2n\] Replacing into the second equation,\[(-2n)-2m-n=0\]\[-2m=3n\]\[m=-\frac{3}{2}n\] Replacing m and l into the first equation, \[3(-2n)-2(-\frac{3}{2}n)+3n=0\]\[-6n+3n+3n=0\] Here's the interesting part,\[n=\mathbb{R} \]n can be any real number, yet l and m depend on the value of n.\[l=-2n\]\[m=-\frac{3}{2}n\]

OpenStudy (souvik):

can it be done by cross multiplication?